OpenStudy (anonymous):

4/7 of a class earned A's 6 did not, how many earned A's? I asked incorrectly last time.

OpenStudy (anonymous):

Lete there are X stidents in the class 4/7 of a class earned A and 6 did not Therefore no. of students who earned= (X-6) Hence: 4/7 of X = X-6 \[\huge \frac{4}{7} \times X = X-6 \rightarrow \huge \frac{4X}{7} = X-6\] \[\rightarrow \huge 4X =7(X-6) \rightarrow 4X = 7X -42\] \[\huge \rightarrow 42= 7X -4X \] \[\huge \rightarrow 3X =42 \rightarrow X = \frac{42}{3} = 14\] Hence 14 students earned A's in the class. @chadadd3

OpenStudy (anonymous):

Sorry, Since X is the total number of students So number of students earned A= X-6= 14-6 = 8 Hence 8 students earned A's. @chadadd3